OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). 7 -14 0 14

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -4}}\quad ,&{\color{blue}{ 8}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ -6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

???

OpenStudy (jdoe0001):

so.... use the given coordinates to get the slope

OpenStudy (anonymous):

-6 - (-4) / -3 - 8 ?

OpenStudy (anonymous):

-2 / -12?

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

-6 - 8/ -3 - 4

OpenStudy (anonymous):

am I on the right track? -4/ -7?

OpenStudy (jdoe0001):

well... one sec

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -4}}\quad ,&{\color{blue}{ 8}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ -6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ -6}}-{\color{blue}{ (-8)}}}{{\color{red}{ -3}}-{\color{red}{ (-4)}}}\implies \cfrac{-6+8}{-3+4}\implies \cfrac{2}{1}\)

OpenStudy (jdoe0001):

hmm wait a sec... I used -8..

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -4}}\quad ,&{\color{blue}{ 8}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ -6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ -6}}-{\color{blue}{ (8)}}}{{\color{red}{ -3}}-{\color{red}{ (-4)}}}\implies \cfrac{-6-8}{-3+4}\implies \cfrac{-14}{1}\)

OpenStudy (anonymous):

ooohhhh

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

I know where I went wrong

OpenStudy (jdoe0001):

- * - = *

OpenStudy (jdoe0001):

hmm - * - = +

OpenStudy (anonymous):

so would -14 be correct?

OpenStudy (jdoe0001):

yeap -14/1 = -14

OpenStudy (jdoe0001):

any number is "any number"/1, the same

OpenStudy (anonymous):

could you help with 1 more?

OpenStudy (jdoe0001):

sure...you can always post anew.. thus if I dunno.... someone else may and we can revise each other

OpenStudy (anonymous):

Alright, follow me to this thread